Hotchkiss Company has two support departments, Maintenance Department (MD) and Personnel Department (PD), and two producing departments, P1 and P2. The Maintenance Department costs of $90,000 are allocated on the basis of standard service hours used. The Personnel Department costs of $9,000 are allocated on the basis of number of employees. The direct costs of Departments P1 and P2 are $36,000 and $60,000, respectively.

Data on standard service hours and number of employees are as follows:

Maint.
Dept.
Person.
Dept.
Dept.
P1
Dept.
P2
Standard service hours used 400 200 1,200   600
Number of employees 20 40   180   180
Direct labor hours 200 200 1,000 1,000


What are the total overhead costs associated with P2 after allocating the Maintenance and Personnel Departments using the direct method?

a.$94,500

b.$90,000

c.$48,000

d.$99,000
Answer #1

Solution:

Maintenance Dept. cost allocated to P2 = $90000 *600/ 1800 = $30,000

Pesonnel dept. cost allocated to P2 = $9000*180/360 = $4,500

Total cost associated with P2 = 60000 +30000 +4500 = $94,500

Hence option "a" is correct.